# Enhanced Muonization by Active-Sterile Neutrino Mixing in Protoneutron Stars

### v1

@@ Abstract
We study $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing in the protoneutron star (PNS) created in a core-collapse supernova (CCSN). We point out the importance of the feedback on the general composition of the PNS in addition to the obvious feedback on the $\nu_{\mu}$ lepton number. We show that for our adopted mixing parameters $\delta m^{2}\sim 10^{2}$ keV2 and $\sin^{2}2\theta$ consistent with the current constraints, sterile neutrino production is dominated by the Mikheyev–Smirnov–Wolfenstein conversion of $\bar{\nu}_{\mu}$ into $\bar{\nu}_{s}$ and that the subsequent escape of $\bar{\nu}_{s}$ increases the $\nu_{\mu}$ lepton number, which in turn enhances muonization of the PNS primarily through $\nu_{\mu}+n\to p+\mu^{-}$. We find that this enhancement is amplified by diffusion of the $\nu_{e}$ and $\nu_{\mu}$ lepton numbers. While these results are qualitatively robust, their quantitative effects on the dynamics and active neutrino emission of CCSNe should be evaluated by including $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing in the simulations.
@@ I Introduction
The rest of the paper is organized as follows. In Section II we discuss the production of sterile neutrinos through $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing in a PNS. Much of the treatment is similar to that of $\nu_{\tau}$-$\nu_{s}$ and $\bar{\nu}_{\tau}$-$\bar{\nu}_{s}$ mixing, and we follow the discussion in Ref. PhysRevD.108.063025 . In Section III we discuss the treatment of the feedback effects of $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing in the presence of muons. In Section IV we present example calculations to illustrate the effects of such mixing on muonization in the PNS. In Section V we discuss our results and give conclusions.
@@ II Production of sterile neutrinos in a PNS
$P_{\rm LZ}=\exp\left(-\frac{\pi\delta m^{2}H_{R}\sin^{2}2\theta}{4E_{R}\cos 2\theta}\right)\,,$ (6)
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ III Feedback from active-sterile neutrino mixing
where $\Gamma_{n\to p}$ and $\Gamma_{p\to n}$ are the net rates for interconverting $n$ and $p$, and $\Gamma_{L_{e}}$ and $\Gamma_{L_{\mu}}$ are the net rates of change in $Y_{e}+Y_{\nu_{e}}$ and $Y_{\mu}+Y_{\nu_{\mu}}$ (see below). The constraints in Eqs. (10a), (10b), (10e), and (10f) correspond to conservation of baryon number, electric charge, and net lepton numbers of the electron and muon types. Because the reactions in Eqs. (9a) and (9b) are very efficient, we have assumed that chemical equilibrium is achieved among $n$, $p$, $e^{\pm}$, $\nu_{e}$, and $\bar{\nu}_{e}$ [Eq. (10c)], and that $Y_{p}$ takes the equilibrium value determined by $\Gamma_{n\to p}$ and $\Gamma_{p\to n}$ [Eq. (10d), see Appendix].
The net rates for interconverting $n$ and $p$ are
$\Gamma_{n\to p}$ $=\Gamma_{\nu_{e}n}+\Gamma_{e^{+}n}+\Gamma_{\nu_{\mu}n}+\Gamma_{\mu^{+}n},$ (11a)
$\Gamma_{p\to n}$ $=\Gamma_{\bar{\nu}_{e}p}+\Gamma_{e^{-}p}+\Gamma_{\bar{\nu}_{\mu}p}+\Gamma_{\mu^{-}p},$ (11b)
where for example, $\Gamma_{\nu_{e}n}$ is the rate per neutron for the forward reaction in Eq. (9a) (see Appendix for details). For the net rates of change in $Y_{e}+Y_{\nu_{e}}$ and $Y_{\mu}+Y_{\nu_{\mu}}$, we consider two scenarios to highlight the effects of sterile neutrino production and illustrate those of active neutrino diffusion. In the first scenario, we ignore the transport of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$ by diffusion and take
@@
In the second scenario, we take
$\Gamma_{L_{e}}$ $=\dot{Y}_{\nu_{e}}^{\rm diff},$ (13a)
$\Gamma_{L_{\mu}}$ $=\dot{Y}_{\nu_{\mu}}^{\rm MSW}+\dot{Y}_{\nu_{\mu}}^{\rm coll}+\dot{Y}_{\nu_{\mu}}^{\rm diff},$ (13b)
where
$\dot{Y}_{\nu_{\beta}}^{\rm diff}=\frac{1}{6\pi G_{F}^{2}n_{b}r^{2}}\frac{\partial}{\partial r}\left(\frac{r^{2}}{n_{b}}\frac{\partial\mu_{\nu_{\beta}}}{\partial r}\right)\,,$ (14)
is the rate of change in $Y_{\nu_{\beta}}$ due to $\nu_{\beta}$ ($\beta=e,\ \mu$) diffusion (e.g., PhysRevD.108.063025 ).
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 3: Evolution of $Y_{e}$, $Y_{\nu_{e}}$, $Y_{\mu}$, and $Y_{\nu_{\mu}}$ as functions of radius $r$ and time $t$ for $m_{s}=7.1$ keV and $\sin^{2}2\theta=7\times 10^{-11}$. The left (right) panels show the results without (with) diffusion of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ IV Results
We now present example calculations of the effects of $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing in a PNS. We expect that in addition to the feedback effects discussed in Section III, such mixing would also affect the evolution of temperature $T$ and density $\rho$. As mentioned in the introduction, such dynamic effects can be properly quantified only by including $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing in CCSN simulations. Our goal here, however, is to elucidate how such mixing enhances muonization. As a reasonable approximation, we limit the duration of our calculations so that the radial profiles of $T$ and $\rho$ can be considered as fixed in time. As the initial conditions, we take a snapshot at time post bounce $t_{\rm pb}=1$ s of a $20\,M_{\odot}$ CCSN model (with the SFHo nuclear equation of state and with muonization) provided by the Garching group 2017PhRvL.119x2702B ; Bollig2016 . Based on the comparison of the conditions at $t_{\rm pb}=1$ s and those at the next output instant $t_{\rm pb}=1.25$ s (see Fig. 2), we consider it reasonable to assume no evolution of $T(r)$ and $\rho(r)$ for a duration of $\Delta t=0.1$ s. We also limit our calculations to within the radius $r=14.7$ km, which corresponds to a temperature of $T=25.6$ MeV and a density of $\rho=8.01\times 10^{13}$ g cm-3, so that all active neutrinos are in the diffusion regime throughout the calculations.
@@
We take $\delta m^{2}\approx m_{s}^{2}$ with $m_{s}=7.1$ keV and $\sin^{2}2\theta=7\times 10^{-11}$, which are suggested by interpreting the X-ray line emission near 3.55 keV from galaxy clusters as due to sterile neutrino decay Bulbul:2014sua ; Boyarsky:2014jta . For comparison, we use $m_{s}=10$ keV and $\sin^{2}2\theta=10^{-12}$, which are chosen for sterile neutrinos to make up all the dark matter Ng:2019gch ; Roach:2022lgo . For convenience, hereafter we refer to the two adopted sets of mixing parameters by the corresponding $m_{s}$ values. Using the rates $\Gamma_{L_{e}}$ and $\Gamma_{L_{\mu}}$ in Eqs. (12a) and (12b) [Eqs. (13a) and (13b)], we show the evolution of $Y_{e}$, $Y_{\nu_{e}}$, $Y_{\mu}$, and $Y_{\nu_{\mu}}$ as functions of radius $r$ and time $t$ in the left (right) panels of Fig. 3 for $m_{s}=7.1$ keV. Similar results for $m_{s}=10$ keV are shown in Fig. 4. The evolution of $Y_{n}$ and $Y_{p}$ (not shown) can be obtained from that of $Y_{e}$ and $Y_{\mu}$ through Eqs. (10a) and (10b).
IV.1 Evolution without diffusion of electron and muon neutrino lepton numbers
@@
We first ignore diffusion of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$ and discuss the results (left panels of Figs. 3 and 4) due to the production and escape of $\nu_{s}$ and $\bar{\nu}_{s}$ only [with $\Gamma_{L_{e}}$ and $\Gamma_{L_{\mu}}$ in Eqs. (12a) and (12b)]. For our adopted mixing parameters, the dominant production of sterile neutrinos is through the MSW conversion of $\bar{\nu}_{\mu}$ into $\bar{\nu}_{s}$. We show the corresponding rate $\dot{Y}_{\nu_{\mu}}^{\rm MSW}$ for change in $Y_{\nu_{\mu}}$ as functions of radius $r$ and time $t$ in the top panels of Fig. 5. Initially $\dot{Y}_{\nu_{\mu}}^{\rm MSW}$ increases with radius as $\bar{\nu}_{e}$ of higher energy goes through MSW resonances at larger radii (bottom panels of Fig. 5). The resulting production and escape of $\bar{\nu}_{s}$ changes $Y_{\nu_{\mu}}$ directly and alter $Y_{\mu}$, $Y_{e}$, $Y_{\nu_{e}}$, $Y_{n}$, and $Y_{p}$ through those charged-current reactions intercoverting $n$ and $p$. Consequently, the potential $V_{\bar{\nu}}$ is changed, which in turn affects the subsequent MSW conversion of $\bar{\nu}_{\mu}$ into $\bar{\nu}_{s}$. These feedback effects are clearly shown in Fig. 5. At the same radius, the resonant energy $E_{R}\approx\delta m^{2}/(2V_{\bar{\nu}})$ increases with time, mostly due to the increase of $Y_{\mu}$ and hence decrease of $V_{\bar{\nu}}\sim\sqrt{2}G_{F}n_{b}[(Y_{n}/2)-Y_{\mu}]$. The evolution of $\dot{Y}_{\nu_{\mu}}^{\rm MSW}$ is more complicated because it depends on both the resonant energy and the evolving energy distribution of $\bar{\nu}_{\mu}$ (due to the change of $\mu_{\nu_{\mu}}$). The general tendency is that $\dot{Y}_{\nu_{\mu}}^{\rm MSW}$ decreases with time, i.e., the feedback of $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing tends to turn off such mixing (e.g., 2011PhRvD..83i3014R ; 2019JCAP…12..019S ; PhysRevD.108.063025 ). The cases of $m_{s}=7.1$ and 10 keV are similar except that the relevant resonant energy is higher (due to a higher $\delta m^{2}$) and the corresponding MSW conversion is less adiabatic (mostly due to a much smaller $\sin^{2}2\theta$) for the latter. The smaller $\dot{Y}_{\nu_{\mu}}^{\rm MSW}$ for $m_{s}=10$ keV leads to smaller effects on $Y_{\nu_{\mu}}$, $Y_{\mu}$, $Y_{e}$, and $Y_{\nu_{e}}$.
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
As shown in the left panels of Figs. 3 and 4, there is little evolution of $Y_{e}$, the evolution of $Y_{\nu_{e}}$ is less than that of $Y_{\nu_{\mu}}$, and the evolution of $Y_{\mu}$ is the largest. The simultaneous changes of $Y_{\mu}$, $Y_{\nu_{\mu}}$, $Y_{\nu_{e}}$, and $Y_{e}$ tend to drive the corresponding particles towards chemical equilibrium. For illustration, we show in the left panel of Fig. 6 the evolution of the chemical potentials $\mu_{e}$, $\mu_{\nu_{e}}$, $\mu_{\mu}$, and $\mu_{\nu_{\mu}}$ for the zone at $r=14.1$ km. Because $(\mu_{e}-\mu_{\nu_{e}})-(\mu_{\mu}-\mu_{\nu_{\mu}})$ stays close to zero, chemical equilibrium among the corresponding particles always holds to good approximation. Note that although we assume $\mu_{\nu_{e}}+\mu_{n}=\mu_{p}+\mu_{e}$ in view of the fast rates of the reactions in Eqs. (9a) and (9b), we do not assume that the rates of the similar reactions involving $\mu^{\pm}$, $\nu_{\mu}$, and $\bar{\nu}_{\mu}$ are fast enough to ensure $\mu_{\nu_{\mu}}+\mu_{n}=\mu_{p}+\mu_{\mu}$. Instead, we assume that $Y_{p}$ takes the equilibrium value determined by the net rates of the reactions interconverting $n$ and $p$, which include the fast reactions in Eqs. (9a) and (9b). Our results show that the charged-current reactions intercoverting $n$ and $p$ by $\mu^{\pm}$, $\nu_{\mu}$, and $\bar{\nu}_{\mu}$ are sufficiently fast to establish $\mu_{\nu_{\mu}}+\mu_{n}\approx\mu_{p}+\mu_{\mu}$, which together with $\mu_{\nu_{e}}+\mu_{n}=\mu_{p}+\mu_{e}$, gives $(\mu_{e}-\mu_{\nu_{e}})-(\mu_{\mu}-\mu_{\nu_{\mu}})\approx 0$.
Refer to captionRefer to caption
Figure 6: Evolution of the chemical potentials $\mu_{e}$, $\mu_{\nu_{e}}$, $\mu_{\mu}$, and $\mu_{\nu_{\mu}}$ as functions of time $t$ for the zone at $r=14.1$ km. The left (right) panels show the results without (with) diffusion of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$ for $m_{s}=7.1$ keV and $\sin^{2}2\theta=7\times 10^{-11}$. Chemical equilibrium among the relevant particles is achieved to good approximation because $(\mu_{e}-\mu_{\nu_{e}})-(\mu_{\mu}-\mu_{\nu_{\mu}})\approx 0$.
The relatively large change of $Y_{\mu}$ compared with much smaller changes of $Y_{\nu_{\mu}}$, $Y_{\nu_{e}}$, and $Y_{e}$ merits discussion. From $(\mu_{e}-\mu_{\nu_{e}})-(\mu_{\mu}-\mu_{\nu_{\mu}})\approx 0$ and the chemical potentials shown in the left panel of Fig. 6, we expect similar changes in $\mu_{\mu}$ and $\mu_{\nu_{\mu}}$. However, for the conditions in the PNS, the increase of $Y_{\mu}$ with $\mu_{\mu}$ is much steeper than that of $Y_{\nu_{\mu}}$ with $\mu_{\nu_{\mu}}$ because of the muon rest mass (see Fig. 7). Therefore, the modest increase of $Y_{\nu_{\mu}}$ mainly due to the MSW conversion of $\bar{\nu}_{\mu}$ into $\bar{\nu}_{s}$ gives rise to the substantial increase of $Y_{\mu}$ through chemical equilibrium.
Refer to captionRefer to caption
IV.2 Evolution with diffusion of electron and muon neutrino lepton numbers
We find that diffusion of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$ enhances the evolution of $Y_{e}$, $Y_{\nu_{e}}$, $Y_{\mu}$, and $Y_{\nu_{\mu}}$ (right panels of Figs. 3 and 4). The rate $\dot{Y}_{\nu_{e}}^{\rm diff}$ for change in $Y_{\nu_{e}}$ due to diffusion is $\lesssim 1$ s-1, and $\dot{Y}_{\nu_{\mu}}^{\rm diff}$ is $\lesssim 3$ s-1. These rates are comparable to $\dot{Y}_{\nu_{\mu}}^{\rm MSW}$ (see Fig. 5). In addition, $\dot{Y}_{\nu_{e}}^{\rm diff}$ and $\dot{Y}_{\nu_{\mu}}^{\rm diff}$ tend to increase with time, in contrast to the decreasing trend of $\dot{Y}_{\nu_{\mu}}^{\rm MSW}$. Consequently, $Y_{\nu_{\mu}}$ reaches significantly higher values with diffusion of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$. Through chemical equilibrium, this increase of $Y_{\nu_{\mu}}$ drives the simultaneous increase of $Y_{\mu}$ similarly to the case without diffusion of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$ (see Fig. 7), except that the change in $\mu_{\mu}$ is approximately the change in $\mu_{\nu_{\mu}}-\mu_{\nu_{e}}$ with significant evolution of $\mu_{\nu_{e}}$ (see the right panel of Fig. 6).

### v2

@@ Abstract
We study $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing in the protoneutron star (PNS) created in a core-collapse supernova (CCSN). We point out the importance of the feedback on the general composition of the PNS in addition to the obvious feedback on the $\nu_{\mu}$ lepton number. We show that for our adopted mixing parameters $\delta m^{2}\sim 10^{2}$ keV2 and $\sin^{2}2\theta$ consistent with the current constraints, sterile neutrino production is dominated by the Mikheyev–Smirnov–Wolfenstein conversion of $\bar{\nu}_{\mu}$ into $\bar{\nu}_{s}$ and that the subsequent escape of $\bar{\nu}_{s}$ increases the $\nu_{\mu}$ lepton number, which in turn enhances muonization of the PNS primarily through $\nu_{\mu}+n\to p+\mu^{-}$. While these results are qualitatively robust, their quantitative effects on the dynamics and active neutrino emission of CCSNe should be evaluated by including $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing in the simulations.
@@ I Introduction
The rest of the paper is organized as follows. In Section II we discuss the production of sterile neutrinos through $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing in a PNS. Much of the treatment is similar to that of $\nu_{\tau}$-$\nu_{s}$ and $\bar{\nu}_{\tau}$-$\bar{\nu}_{s}$ mixing, and we closely follow the discussion in Ref. PhysRevD.108.063025 . In Section III we discuss the treatment of the feedback effects of $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing in the presence of muons. In Section IV we present example calculations to illustrate the effects of such mixing on muonization in the PNS. In Section V we discuss our results and give conclusions.
@@ II Production of sterile neutrinos in a PNS
$P_{\rm LZ}=\exp\left(-\frac{\pi\delta m^{2}H_{R}\sin^{2}2\theta}{4E_{R}\cos 2\theta}\right)\,$ (6)
@@ III Feedback from active-sterile neutrino mixing
where $\Gamma_{L_{e}}$ and $\Gamma_{L_{\mu}}$ are the net rates of change in $Y_{e}+Y_{\nu_{e}}$ and $Y_{\mu}+Y_{\nu_{\mu}}$ (see below). The constraints in Eqs. (10a), (10b), (10e), and (10f) correspond to conservation of baryon number, electric charge, and net lepton numbers of the electron and muon types. We have assumed that chemical equilibrium is achieved among $n$, $p$, $e^{\pm}$, $\nu_{e}$, and $\bar{\nu}_{e}$ [Eq. (10c)], and also among $n$, $p$, $\mu^{\pm}$, $\nu_{\mu}$, and $\bar{\nu}_{\mu}$ [Eq. (10d)]. This assumption is consistent with the Fermi-Dirac energy distributions adopted for the above particles and with the fast rates (see Appendix) for interconverting $n$ and $p$ by these particles.
For the net rates of change in $Y_{e}+Y_{\nu_{e}}$ and $Y_{\mu}+Y_{\nu_{\mu}}$, we ignore the transport of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$ by diffusion and take
@@
The above approximation highlights the effects of sterile neutrino production and its validity will be discussed in Section V.
Figure 3: Evolution of $Y_{e}$, $Y_{\nu_{e}}$, $Y_{\mu}$, and $Y_{\nu_{\mu}}$ as functions of radius $r$ and time $t$ for $m_{s}=7.1$ keV and $\sin^{2}2\theta=7\times 10^{-11}$. Diffusion of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$ is ignored.
@@ IV Results
We now present example calculations of the effects of $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing in a PNS. We expect that in addition to the feedback effects discussed in Section III, such mixing would also affect the evolution of temperature $T$ and density $\rho$. As mentioned in the introduction, such dynamic effects can be properly quantified only by including $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing in CCSN simulations. Our goal here, however, is to elucidate how such mixing enhances muonization. As a reasonable approximation, we limit the duration of our calculations so that the radial profiles of $T$ and $\rho$ can be considered as fixed in time. As the initial conditions, we take a snapshot at time post bounce $t_{\rm pb}=1$ s of a $20\,M_{\odot}$ CCSN model (with the SFHo nuclear equation of state and with muonization) provided by the Garching group 2017PhRvL.119x2702B ; Bollig2016 . Based on the comparison of the conditions at $t_{\rm pb}=1$ s and those at the next output instant $t_{\rm pb}=1.25$ s (see Fig. 2), we consider it reasonable to assume no evolution of $T(r)$ and $\rho(r)$ for a duration of $\Delta t=0.1$ s. We also limit our calculations to within the radius $r=14.7$ km, which corresponds to a temperature of $T=25.6$ MeV and a density of $\rho=8.01\times 10^{13}$ g cm-3, so that all active neutrinos are in the trapped regime throughout the calculations.
@@
We take $\delta m^{2}\approx m_{s}^{2}$ with $m_{s}=7.1$ keV and $\sin^{2}2\theta=7\times 10^{-11}$, which are suggested by interpreting the X-ray line emission near 3.55 keV from galaxy clusters as due to sterile neutrino decay Bulbul:2014sua ; Boyarsky:2014jta . For comparison, we use $m_{s}=10$ keV and $\sin^{2}2\theta=10^{-12}$, which are chosen for sterile neutrinos to make up all the dark matter Ng:2019gch ; Roach:2022lgo . For convenience, hereafter we refer to the two adopted sets of mixing parameters by the corresponding $m_{s}$ values. Using the rates $\Gamma_{L_{e}}$ and $\Gamma_{L_{\mu}}$ in Eqs. (11a) and (11b), we show the evolution of $Y_{e}$ (top left), $Y_{\nu_{e}}$ (top right), $Y_{\mu}$ (bottom left), and $Y_{\nu_{\mu}}$ (bottom right) as functions of radius $r$ and time $t$ in Figs. 3 and 4 for $m_{s}=7.1$ and 10 keV, respectively. The evolution of $Y_{n}$ and $Y_{p}$ (not shown) can be obtained from that of $Y_{e}$ and $Y_{\mu}$ through Eqs. (10a) and (10b).
As shown in Figs. 3 and 4, there is little evolution of $Y_{e}$, the evolution of $Y_{\nu_{e}}$ is less than that of $Y_{\nu_{\mu}}$, and the evolution of $Y_{\mu}$ is the largest. This comparison is reflected by changes of the corresponding chemical potentials. For illustration, we show in the left (right) panel of Fig. 5 the evolution of the chemical potentials $\mu_{e}$, $\mu_{\nu_{e}}$, $\mu_{\mu}$, and $\mu_{\nu_{\mu}}$ for the zone at $r=14.1$ km for $m_{s}=7.1$ (10) keV. It can be seen that the changes $|\Delta\mu_{e}|$ and $|\Delta\mu_{\nu_{e}}|$ are much smaller than $|\Delta\mu_{\nu_{\mu}}|$ and $|\Delta\mu_{\mu}|$, which along with $(\mu_{e}-\mu_{\nu_{e}})-(\mu_{\mu}-\mu_{\nu_{\mu}})=0$ for chemical equilibrium [see Eqs. (10c) and (10d)], gives $\Delta\mu_{\mu}\sim\Delta\mu_{\nu_{\mu}}$. For the conditions in the PNS, the increase of $Y_{\mu}$ with $\mu_{\mu}$ is much steeper than that of $Y_{\nu_{\mu}}$ with $\mu_{\nu_{\mu}}$ because of the muon rest mass (see Fig. 6). Therefore, the modest increase of $Y_{\nu_{\mu}}$ mainly due to the MSW conversion of $\bar{\nu}_{\mu}$ into $\bar{\nu}_{s}$ gives rise to the substantial increase of $Y_{\mu}$ through chemical equilibrium.
Figure 5: Evolution of the chemical potentials $\mu_{e}$, $\mu_{\nu_{e}}$, $\mu_{\mu}$, and $\mu_{\nu_{\mu}}$ as functions of time $t$ for the zone at $r=14.1$ km. The left (right) panel shows the results for $m_{s}=7.1\,(10)$ keV and $\sin^{2}2\theta=7\times 10^{-11}\,(10^{-12})$. Chemical equilibrium among the relevant particles gives $(\mu_{e}-\mu_{\nu_{e}})-(\mu_{\mu}-\mu_{\nu_{\mu}})=0$.
@@
To understand better the dominant production of sterile neutrinos through the MSW conversion of $\bar{\nu}_{\mu}$ into $\bar{\nu}_{s}$, we show the corresponding rate $\dot{Y}_{\nu_{\mu}}^{\rm MSW}$ for change in $Y_{\nu_{\mu}}$ as functions of radius $r$ and time $t$ in the top panels of Fig. 7. Initially $\dot{Y}_{\nu_{\mu}}^{\rm MSW}$ increases with radius as $\bar{\nu}_{\mu}$ of higher energy go through MSW resonances at larger radii (bottom panels of Fig. 7). The resulting production and escape of $\bar{\nu}_{s}$ change $Y_{\nu_{\mu}}$ directly and alter $Y_{\mu}$, $Y_{e}$, $Y_{\nu_{e}}$, $Y_{n}$, and $Y_{p}$ through chemical equilibrium. Consequently, the potential $V_{\bar{\nu}}$ is changed, which in turn affects the subsequent MSW conversion of $\bar{\nu}_{\mu}$ into $\bar{\nu}_{s}$. These feedback effects are clearly shown in Fig. 7. At the same radius, the resonant energy $E_{R}\approx\delta m^{2}/(2V_{\bar{\nu}})$ increases with time, mostly due to the increase of $Y_{\mu}$ and hence decrease of $V_{\bar{\nu}}\sim\sqrt{2}G_{F}n_{b}[(Y_{n}/2)-Y_{\mu}]$. The evolution of $\dot{Y}_{\nu_{\mu}}^{\rm MSW}$ is more complicated because it depends on both the resonant energy and the evolving energy distribution of $\bar{\nu}_{\mu}$ (due to the change of $\mu_{\nu_{\mu}}$). The general tendency is that $\dot{Y}_{\nu_{\mu}}^{\rm MSW}$ decreases with time, i.e., the feedback of $\nu_{\mu}$-$\nu_{s}$ and $\bar{\nu}_{\mu}$-$\bar{\nu}_{s}$ mixing tends to turn off such mixing (e.g., 2011PhRvD..83i3014R ; 2019JCAP…12..019S ; PhysRevD.108.063025 ). While the cases of $m_{s}=7.1$ and 10 keV are similar, the relevant resonant energy is higher (due to a higher $\delta m^{2}$) and the corresponding MSW conversion is less adiabatic (mostly due to a much smaller $\sin^{2}2\theta$) for the latter. Consequently, the evolution of $Y_{\mu}$ is significantly different for the two cases (the differences in the evolution of $Y_{\nu_{e}}$ and $Y_{\nu_{\mu}}$ are also noticeable, see Figs. 3 and 4).